A cable exerts a constant upward tension of magnitude 2.5 x 104 N on a fully loaded elevator with a mass of 1.8 x 103 kg as it rises to a vertical distance of 25.5 m.


1
Expert's answer
2021-01-27T15:26:58-0500

a) Let's calculate the work done by the tension force on the elevator:


"W_T=F_Td=2.5\cdot10^4\ N\cdot25.5\ m=6.37\cdot10^5\ J."

b) Let's calculate the work done by the force of gravity on the elevator:


"W_g=F_gd=mgd=1.8\cdot10^3\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot25.5\ m=4.5\cdot10^5\ J."